# Integrated GARCH-GRU in Financial Volatility Forecasting

### v1

@@ 4.2 Model Evaluation
Table.3 presents the empirical forecasting results for all models, where performance metrics are reported as averages over 20 runs with different random seeds. The exception is the traditional GARCH models, which are deterministic and do not involve stochastic training procedures. The table compares the forecasting accuracy of the proposed integrated GARCH-GRU model against a comprehensive set of benchmarks, including various hybrid model variants, standalone deep learning architectures, and traditional econometric models. Here the proposed integrated GARCH-GRU model, denoted as ‘eh_GARCH_GRU’, represents an enhanced hybrid architecture that unifies GARCH dynamics with the GRU structure. This comparative framework allows for a robust assessment of the GARCH-GRU model’s effectiveness across different forecasting horizons and datasets. In addition, the in-sample estimation results for the GARCH component parameters across all relevant models are presented in Appendix.A, providing further insights into the statistical properties and internal dynamics of the hybrid architectures.
Table 3: Model Performance Comparison
Model eh_GARCH_GRU eh_GARCH_LSTM bl_GARCH_LSTM pl_GARCH_LSTM GRU LSTM Transformer GARCH(1,1) GJR-GARCH
Risk Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
S&P 500 1D 0.0202 (0.0012) 0.0976 (0.0047) 0.0256 (0.0020) 0.1073 (0.0055) 0.0616 (0.0129) 0.1940 (0.0236) 0.0224 (0.0024) 0.1006 (0.0067) 0.0217 (0.0014) 0.0989 (0.0046) 0.0230 (0.0022) 0.1026 (0.0056) 0.0234 (0.0062) 0.1040 (0.0164) 0.0795 0.2174 0.0863 0.2124
3D 0.0375 (0.0013) 0.1383 (0.0059) 0.0483 (0.0042) 0.1590 (0.0075) 0.0730 (0.0061) 0.2064 (0.0112) 0.0409 (0.0041) 0.1432 (0.0084) 0.0407 (0.0052) 0.1454 (0.0129) 0.0414 (0.0052) 0.1450 (0.0097) 0.0493 (0.0041) 0.1556 (0.0120) 0.1053 0.2529 0.1138 0.2519
1W 0.0785 (0.0035) 0.2077 (0.0100) 0.1180 (0.0162) 0.2534 (0.0166) 0.0822 (0.0121) 0.2177 (0.0148) 0.0896 (0.0269) 0.2159 (0.0156) 0.0983 (0.0109) 0.2302 (0.0173) 0.0802 (0.0048) 0.2132 (0.0085) 0.0799 (0.0025) 0.2082 (0.0075) 0.1635 0.3125 0.1961 0.3322
DJI 1D 0.0225 (0.0013) 0.1036 (0.0064) 0.0252 (0.0017) 0.1106 (0.0050) 0.0596 (0.0135) 0.1981 (0.0215) 0.0216 (0.0024 ) 0.1022 (0.0101) 0.0206 (0.0013) 0.1017 (0.0064) 0.0224 (0.0015) 0.1036 (0.0066) 0.0309 (0.0042) 0.1228 (0.0101) 0.0767 0.2101 0.0925 0.2250
3D 0.0360 (0.0016) 0.1388 (0.0045) 0.0490 (0.0041) 0.1591 (0.0058) 0.0676 (0.0046) 0.2001 (0.0084) 0.0376 (0.0024) 0.1389 (0.0040) 0.0333 (0.0009) 0.1330 (0.0013) 0.0385 (0.0041) 0.1406 (0.0080) 0.0520 (0.0086) 0.1650 (0.0105) 0.0979 0.2396 0.1202 0.2601
1W 0.0759 (0.0030) 0.2040 (0.0072) 0.1090 (0.0157) 0.2388 (0.0146) 0.7463 (0.4080) 0.5211 (0.1607) 0.0836 (0.0082) 0.2099 (0.0122) 0.0931 (0.0093) 0.2276 (0.0117) 0.0847 (0.0115) 0.2121 (0.0149) 0.0860 (0.0071) 0.2184 (0.0114) 0.1470 0.2906 0.1917 0.3257
NASDAQ 1D 0.0288 (0.0015) 0.1179 (0.0056) 0.0369 (0.0028) 0.1317 (0.0062) 0.4576 (1.5881) 0.3612 (0.5266) 0.0372 (0.0029) 0.1286 (0.0043) 0.0376 (0.0881) 0.1323 (0.0713) 0.0328 (0.0030) 0.1217 (0.0066) 0.0337 (0.0045) 0.1253 (0.0129) 0.1342 0.2761 0.1446 0.2766
3D 0.0572 (0.0017) 0.1717 (0.0053) 0.0670 (0.0055) 0.1869 (0.0091) 0.0723 (0.0045) 0.2040 (0.0084) 0.0698 (0.0041) 0.1909 (0.0076) 0.0695 (0.0056) 0.1879 (0.0095) 0.0607 (0.0066) 0.1764 (0.0112) 0.0780 (0.0090) 0.1984 (0.0119) 0.1659 0.3213 0.1835 0.3233
1W 0.1086 (0.0027) 0.2476 (0.0047) 0.1542 (0.0182) 0.2919 (0.0178) 0.1248 (0.0279) 0.2838 (0.0802) 0.1202 (0.0071) 0.2597 (0.0095) 0.1344 (0.0207) 0.2721 (0.0187) 0.1353 (0.0148) 0.2720 (0.0153) 0.1386 (0.0149) 0.2785 (0.0198) 0.2594 0.4017 0.3136 0.4169
‘eh_GARCH_GRU’ and ‘eh_GARCH_LSTM’ denote the proposed end-to-end integrated models. ‘bl_GARCH_LSTM’ refers to the embedded integration approach that modifies the output gate within the LSTM architecture. ‘pl_GARCH_LSTM’ and ‘pl_GARCH_GRU’ represent pipeline-based integration models, where GARCH and neural components are trained sequentially. Values in parentheses indicate standard deviations computed from 20 independent runs with different random seeds. As the GARCH(1,1) and GJR-GARCH models are deterministic and do not involve stochastic training procedures, standard deviations are not reported for these benchmarks.
In the overall comparison of the model forecasting errors across the three datasets, our integrated GARCH-GRU model consistently outperforms other benchmark models, achieving the lowest MSE and MAE values in most forecasting scenarios. This superior performance is especially evident in medium- and long-term forecasting horizons (3-day and 1-week) across the S&P 500 and NASDAQ indices. An exception arises in the case of the Dow Jones Industrial Average (DJI) at the 1-day and 3-day forecast horizons, where the GARCH-GRU model exhibits slightly higher forecasting errors than the baseline GRU model. This performance difference can be attributed to several factors. Notably, the DJI consists of a smaller and more stable group of blue-chip stocks, which typically exhibit lower short-term volatility and weaker clustering effects compared to the broader S&P 500 or the more volatility-prone, technology-driven NASDAQ. In such settings, the benefits of incorporating GARCH-style persistence into the model may be limited, and purely data-driven architectures like GRU may offer better adaptability to short-term fluctuations that are less driven by persistent volatility shocks. Nonetheless, the differences in errors are relatively small and statistically insignificant, and the integrated GARCH-GRU model still provides a more interpretable and principled framework.
When compared to the integrated GARCH-LSTM model, the GARCH-GRU consistently outperforms across all datasets and forecast horizons. On the S&P 500, for example, the GARCH-GRU model achieves lower errors at every horizon, with the largest gap observed at the 1-week forecast, where the MSE is reduced from 0.1180 to 0.0785, and MAE is reduced from 0.2534 to 0.2077. Similar improvements are seen on the NASDAQ and Dow Jones (DJI) indices. This performance gap reflects the efficiency of the GRU architecture, which, despite having a more compact gating structure, proves more effective when tightly integrated with GARCH dynamics. The GRU’s reduced parameter count may also contribute to better generalization in volatility forecasting tasks, especially under data constraints or high volatility regimes.
Beyond its direct comparison with the GARCH-LSTM model, the proposed GARCH-GRU architecture also outperforms alternative hybrid configurations, including the embedded GARCH-LSTM (bl_GARCH_LSTM), which replaces the LSTM output gate with a GARCH regression, and the pipeline-based GARCH-LSTM (pl_GARCH_LSTM), which sequentially combines GARCH and neural network components. These alternative designs often suffer from limited integration or strong constraints, thereby constraining the model’s ability to fully capture the complex interplay between temporal dynamics and volatility behavior. For instance, on the S&P 500 index at the 1-day horizon, the pipeline model (pl_GARCH_LSTM) yields an MSE of 0.0224, while the embedded model (bl_GARCH_LSTM) performs significantly worse with an MSE of 0.0616. In contrast, the GARCH-GRU achieves a lower MSE of 0.0202, demonstrating its superior forecasting precision. These results highlight the advantages of a tightly coupled, end-to-end architecture, where GARCH dynamics are structurally embedded within the GRU’s update equations. Such integration facilitates more expressive representation learning and enables the model to coherently internalize both nonlinear temporal dependencies and heteroskedastic variance patterns, leading to improved volatility forecasts.
The proposed GARCH-GRU model also demonstrates clear advantages over standalone neural network architectures, including standard GRU, LSTM, and Transformer models. Although these models are capable of learning nonlinear temporal patterns, they lack explicit mechanisms for modeling volatility ‘stylized facts’ such as clustering or long memory. The GARCH-GRU model bridges this gap by embedding statistical volatility dynamics into the hidden state transitions. On the Dow Jones (DJI) index at the 1-week horizon, the standalone GRU and LSTM models yield MSEs of 0.0931 and 0.0847, respectively, while the GARCH-GRU model reduces the error to 0.0759. While the Transformer model performs competitively on certain short-term horizons, its performance degrades over longer horizon, suggesting that its attention mechanism lacks the necessary structural bias for volatility modeling. These results underline the value of combining econometric priors with neural architectures to improve robustness and long-horizon predictive power.
In comparison with the traditional econometric models, such as GARCH(1,1) and GJR-GARCH, the GARCH-GRU model achieves substantial improvements across all datasets and forecasting horizons. The classical models, while interpretable and grounded in financial theory, are limited in their capacity to capture nonlinear dependencies and complex temporal structures. Across all datasets and horizons, the GARCH-GRU model achieves large reductions in both MSE and MAE. For example, on the S&P 500 dataset for 1-day forecasts, GARCH(1,1) and GJR-GARCH models report MSEs of 0.0795 and 0.0863, respectively, compared to 0.0202 from the proposed model, which reaches a reduction of over 70%. Similar margins can be seen in longer-horizon forecasts and other datasets, underscoring the limitations of purely statistical models in modern financial environments characterized by nonlinear and regime-dependent volatility dynamics.
Furthermore, to assess computational efficiency, we measured the average training time for both the integrated GARCH-GRU and GARCH-LSTM models across 20 runs with different random seeds. The results show that the GARCH-GRU model requires only 133.52 seconds on average, compared to 351.57 seconds for the GARCH-LSTM model. It indicates that the GARCH-GRU model achieves approximately 62% reduction in training time. This substantial reduction in training time highlights the computational efficiency of the GRU architecture, which benefits from a simpler gating mechanism while preserving strong sequence modeling capabilities. The ability of the GARCH-GRU model to achieve superior forecasting accuracy with significantly lower computational cost further underscores the advantage of integrating GARCH dynamics with GRU. This makes the proposed model particularly appealing for large-scale or real-time financial forecasting applications, where both performance and efficiency are critical.

### v2

@@ 4.2 Model Evaluation
4.2.1 Normal-Period Evaluation (2019 Out-of-Sample)
Having established the experimental setup, we first examine the empirical forecasting performance of the proposed GARCH-RNN models during the 2019 out-of-sample period. Tables.23 report the mean accuracy metrics across 20 independent runs for all models, with deterministic GARCH-type benchmarks shown without standard deviations. By comparing the proposed integrated GARCH-GRU and GARCH-LSTM structures against classical econometric models, standard deep learning architectures, and pipeline-style hybrids, the evaluation provides a rigorous and comprehensive assessment of forecasting capability.
Across all three indices, the results consistently show that the GARCH-RNN (both GARCH-GRU and GARCH-LSTM) framework outperforms all benchmark models. This is most clearly reflected in the mean squared error (MSE) shown in Table.2, which places greater weight on large deviations and therefore serves as a stringent measure of forecasting precision. The GARCH-GRU attains the lowest MSEs for the S&P 500 at all forecasting horizons, illustrating its strong ability to model short-term volatility shocks while preserving the persistence characteristic of financial volatility. Its advantage becomes even more pronounced in the NASDAQ dataset at medium and long horizons (3- and 7-day), where the GARCH-GRU clearly dominates. The Transformer performs competitively at very short horizons, but its accuracy deteriorates more rapidly as the horizon lengthens, whereas the GARCH-GRU maintains clear stability. For the Dow Jones (DJI) dataset, the GARCH-LSTM delivers a slightly lower MSE at the 1-day and 3-day horizon, indicating its strength in capturing highly localized short-term patterns. However, this advantage disappears as the forecast window extends. At the 7-day horizon, although the Transformer performs slightly better, both GARCH-GRU and GARCH-LSTM remain competitive, with the GARCH-GRU outperforming the GARCH-LSTM. These results suggest that although the LSTM-based variant can excel in very short-horizon settings for certain markets, the GRU-based architecture offers stronger robustness, especially when multi-step temporal dynamics are involved. The pipeline hybrid (bm_GARCH_LSTM), which relies on sequential rather than joint learning, performs markedly worse, confirming that the performance gains arise from embedding GARCH dynamics directly into the recurrent architecture rather than merely appending volatility features to a standard neural network. Traditional GARCH and GJR-GARCH models yield substantially higher errors across all datasets, demonstrating the limitations of fixed parametric forms in capturing the nonlinear and regime-dependent behavior of realized volatility.
Table 2: Forecasting Performance in Normal-Period (Metric: MSE)
Dataset Horizon GARCH_GRU GARCH_LSTM bm_GARCH_LSTM Transformer GARCH(1,1) GJR-GARCH
S&P 500 1
0.0183
0.0223
0.0539
0.0191
0.0795 0.0863
3
0.0360
0.0374
0.0720
0.0431
0.1053 0.1138
7
0.0769
0.0794
0.1121
0.0801
0.1635 0.1961
DJI 1
0.0236
0.0188
0.0564
0.0209
0.0767 0.0925
3
0.0392
0.0380
0.0691
0.0445
0.0979 0.1202
7
0.0834
0.0866
0.0988
0.0782
0.1470 0.1917
NASDAQ 1
0.0353
0.0324
0.0841
0.0308
0.1342 0.1446
3
0.0604
0.0624
0.1116
0.0652
0.1659 0.1835
7
0.1179
0.1179
0.1527
0.1192
0.2594 0.3136
Entries are mean (standard deviation) over 20 seeds where applicable. Deterministic GARCH benchmarks omit standard deviations. ‘bm_GARCH_LSTM’ denotes the benchmark pipeline-based hybrid model, where the GARCH component and the LSTM network are trained sequentially rather than jointly.
The out-of-sample $R^{2}$ results in Table.3 reinforce these observations. This metric evaluates how well the model’s predictions match the observed test data on average, providing an interpretable measure of explanatory power in an out-of-sample context. For all datasets, the GARCH-RNN, especially GARCH-GRU, yields the highest or near-highest $R^{2}$ values, suggesting that it captures a larger proportion of the variance in realized volatility compared to competing approaches. The advantage is especially pronounced for the S&P 500, where the $R^{2}$ remains above 0.87 at horizon 1 and above 0.49 even at horizon 7, demonstrating strong explanatory power and temporal stability. The Transformer models also exhibit solid performance, though the GARCH-RNN’s ability to retain information over longer sequences translates into slightly higher predictability for extended horizons. In contrast, the traditional GARCH and GJR-GARCH models show severely degraded $R^{2}$ values, underscoring their limited capacity to generalize beyond short-term conditional variance estimation. This gap highlights the necessity of neural augmentations for effectively modeling high-dimensional and regime-dependent volatility dynamics.
Table 3: Forecasting Performance in Normal-Period (Metric: $R^{2}$)
Dataset Horizon GARCH_GRU GARCH_LSTM bm_GARCH_LSTM Transformer GARCH(1,1) GJR-GARCH
S&P 500 1
0.8793
0.8532
0.6448
0.8744
-415.9270 -517.5718
3
0.7618
0.7527
0.5242
0.7148
0.3049 0.2493
7
0.4900
0.4731
0.2557
0.4686
-0.0843 -0.3007
DJI 1
0.8285
0.8630
0.5894
0.8483
-412.3757 -519.1085
3
0.7140
0.7225
0.4956
0.6755
0.2865 0.1243
7
0.3877
0.3643
0.2740
0.4262
-0.0784 -0.4064
NASDAQ 1
0.8347
0.8484
0.6067
0.8559
-433.5734 -520.2138
3
0.7176
0.7083
0.4780
0.6950
0.2232 0.1407
7
0.4457
0.4458
0.2817
0.4399
-0.2197 -0.4744
Entries are mean (standard deviation) over 20 seeds where applicable. Deterministic GARCH benchmarks omit standard deviations. ‘bm_GARCH_LSTM’ denotes the pipeline-based hybrid model. Negative OOS $R^{2}$ values naturally arise under a test-set–mean baseline and simply indicate poorer predictive fit than the mean benchmark.
The MAE and SMAPE results for the evaluation in 2019 are reported in Appendix.1011 as complementary evidence. These two metrics measure absolute and proportional forecast deviations and therefore, provide an additional perspective alongside the variance-based measures emphasized in the main text. As shown in the appendix tables, the GARCH-RNN models remain competitive under both metrics, although the Transformer occasionally achieves slightly lower errors due to its ability to capture very local patterns in a relatively stable volatility environment. In normal-market conditions, however, variance-based measures such as MSE and out-of-sample $R^{2}$ are more informative and discriminative, which motivates their use as the primary metrics for the analysis in 2019. By contrast, during the 2020 stress period (Covid-pandemic), characterized by large, abrupt, and highly heterogeneous volatility movements, MAE and SMAPE provide more reliable assessments of absolute and relative forecast accuracy. Accordingly, the subsequent stress-period analysis places greater emphasis on these two metrics.
(a) bm_GARCH_LSTM vs GARCH_RNN
(b) Transformer vs GARCH_RNN
(c) GARCH vs GARCH_RNN
(d) GJR–GARCH vs GARCH_RNN
Figure 3: Response comparison to realized volatility. Each panel overlays realized volatility (black) with the GARCH_GRU forecast (blue), GARCH_LSTM forecast (orange) and a comparator model (green).
To complement the quantitative accuracy metrics, we also examine the time-series behavior of the forecasts. Figure.3 plots the predicted volatility paths of GARCH-GRU, GARCH-LSTM, and all benchmark models against realized volatility over the 2019 test window. Both integrated GARCH-RNN models visually track the realized series remarkably well, with close alignment to the timing and magnitude of major volatility swings. This qualitative evidence reinforces the earlier numerical results: the embedded GARCH dynamics allow both recurrent architectures to adapt quickly to changing market conditions. Across all four panels in Figure.3, the GARCH-GRU (blue) in particular aligns more promptly with abrupt movements in realized volatility (black). During the early-August volatility surge, GARCH-GRU rises to the new level within one to two days and subsequently relaxes almost immediately as volatility recedes. GARCH-LSTM (orange) exhibits a similarly strong fit, though with slightly smoother transitions that occasionally delay its turning-point response. In contrast, the classical GARCH and GJR-GARCH models (green in panels 3(c)3(d)) display slower mean-reverting adjustments and frequent overshooting, consistent with their rigid parametric structure. Relative to neural baselines, the difference is equally notable. The two-stage hybrid (bm_GARCH_LSTM) shows visibly noisier and less synchronized movements, failing to capture the sharp August and October spikes. The Transformer baseline tracks medium-range fluctuations reasonably well but exhibits noticeable lag at several turning points, such as the late-September trough and early-October jump, where both GARCH-GRU and GARCH-LSTM adjust more quickly and with fewer oscillations. These visual comparisons corroborate that the advantage of the proposed models stems not only from lower error statistics but also from their ability to respond dynamically and coherently to regime shifts in volatility.
Table 4: Volatility Forecasting Performance High-Volatility Subsample of Normal-Period($q=0.9$)
Model MSE MAE SMAPE
GARCH_GRU 0.0679 0.1982 0.1343
GARCH-LSTM 0.1056 0.2380 0.1589
bm_GARCH-LSTM 0.1355 0.2958 0.2127
Transformer 0.0753 0.2027 0.1398
GARCH 0.1896 0.3524 0.2661
GJR-GARCH 0.2472 0.3829 0.2447
Results are based on 1-day-ahead forecasts using the S&P 500 in the 2019 normal-period test sample.
Additional perspective is provided by Table.4, which evaluates model performance within a high-volatility subsample (defined by the top 10% of realized volatility observations). This setting mimics market stress periods, where accurate volatility prediction becomes most relevant for financial stability and risk control. The GARCH-GRU remains the strongest performer across all three metrics (MSE, MAE, and SMAPE), indicating that it not only captures the magnitude of volatility spikes more accurately but also adapts more rapidly to the abrupt, nonlinear fluctuations characteristic of turbulent markets.
For completeness, we also compare the unified GARCH–GRU/LSTM (GARCH-RNN) with standard GRU/LSTM model trained under the same settings. While the GARCH-RNN does not consistently outperform the plain NN across all datasets and horizons, their performance differences are generally small. This indicates that incorporating the GARCH mechanism preserves the NN’s predictive capability while providing stronger theoretical interpretability and stability across different market conditions.
Table 5: Estimated GARCH Parameters for S&P 500
Model Horizon $\boldsymbol{\omega}$ $\boldsymbol{\alpha}$ $\boldsymbol{\beta}$ $\boldsymbol{\alpha+\beta}$
GARCH_GRU $h=1$ 0.0244 0.1063 0.8637 0.9700
$h=3$ 0.0300 0.1102 0.8598 0.9700
$h=7$ 0.0223 0.1106 0.8592 0.9689
GARCH_LSTM $h=1$ 0.0260 0.0861 0.8839 0.9700
$h=3$ 0.0214 0.0863 0.8837 0.9700
$h=7$ 0.0205 0.0863 0.8836 0.9699
GARCH(1,1) $h=1$ 0.0940 0.1328 0.7125 0.8453
$h=3$ 0.0969 0.1317 0.7135 0.8452
$h=7$ 0.1003 0.1351 0.7188 0.8539
GJR-GARCH $h=1$ 0.1069 0.1135 0.5660 0.6795
$h=3$ 0.1139 0.1135 0.5611 0.6746
$h=7$ 0.1275 0.1135 0.5550 0.6685
For GJR-GARCH, the leverage parameter $\gamma$ is estimated but omitted here for comparability.
A further advantage of the proposed GARCH-RNN architecture lies in its ability to preserve interpretable GARCH dynamics while jointly learning nonlinear recurrent structure, as reflected in Eq.21. Table.5 reports the estimated $\omega$, $\alpha$, and $\beta$ coefficients extracted from the embedded GARCH layer for the S&P 500. Both GARCH-GRU and GARCH-LSTM produce economically plausible coefficients. The shock response $\alpha$ remains moderate ( $\approx 0.09-0.11$ ), the persistence parameter $\beta$ remains high $(\approx 0.86-0.88)$, and the resulting $\alpha+\beta\approx 0.97$ closely matches well-documented persistence in equity volatility. Notably, these parameter values remain stable across forecasting horizons, indicating that unified training does not distort the underlying GARCH dynamics but instead complements them with nonlinear recurrent structure. Compared with classical GARCH and GJR-GARCH estimates, the embedded GARCH layers yield lower unconditional variance ($\omega$) and substantially higher persistence, consistent with the smoother and more adaptive volatility transitions observed in the time-series plots. These results demonstrate that the GARCH-RNN models retain interpretable GARCH behavior while benefiting from neural gating mechanisms, helping explain their superior predictive accuracy across the evaluated horizons and datasets.
4.2.2 Stress-Period Evaluation (2020 COVID Pandemic)
Having established model performance under normal market conditions, we now examine the behavior of all models during the 2020 COVID-19 period, a regime characterized by extreme volatility, rapid structural breaks, and unusually persistent shocks. This period provides a stringent test of robustness, as forecasting errors typically amplify when volatility levels rise sharply and shift unpredictably. In this setting, absolute- and scale-adjusted deviation measures such as MAE and SMAPE become particularly informative, since they directly quantify the magnitude and proportional severity of prediction errors during large fluctuations.
Table 6: Forecasting Performance in Stress-Period (Metric: MAE)
Dataset Horizon GARCH_GRU GARCH_LSTM bm_GARCH_LSTM Transformer GARCH(1,1) GJR-GARCH
S&P 500 1
0.3408
0.3360
0.4255
0.4358
0.4441 0.4764
3
0.4451
0.4632
0.5048
0.5293
0.5290 0.5552
7
0.6361
0.6358
0.7011
0.6622
0.6691 0.7151
DJI 1
0.4581
0.3932
0.4981
0.5255
0.4570 0.4842
3
0.5295
0.5121
0.5613
0.6268
0.5507 0.5682
7
0.7054
0.7507
0.7298
0.7446
0.6945 0.7443
NASDAQ 1
0.3939
0.3682
0.4833
0.4375
0.4857 0.4911
3
0.5013
0.4781
0.5313
0.5424
0.5827 0.5791
7
0.6955
0.6703
0.7029
0.6868
0.7387 0.7568
Entries are Stress-period (2020) MAE mean (standard deviation) across 20 seeds where applicable. Deterministic GARCH benchmarks omit standard deviations.
Table 7: Forecasting Performance in Stress-Period (Metric: SMAPE)
Dataset Horizon GARCH_GRU GARCH_LSTM bm_GARCH_LSTM Transformer GARCH(1,1) GJR-GARCH
S&P 500 1
0.1804
0.1807
0.2900
0.2221
0.7180 0.7268
3
0.2482
0.2515
0.3229
0.2872
0.3401 0.3430
7
0.3649
0.3606
0.4186
0.3732
0.3993 0.4156
DJI 1
0.2070
0.1850
0.2867
0.2415
0.7056 0.7080
3
0.2573
0.2553
0.3153
0.3059
0.3171 0.3307
7
0.3575
0.3772
0.3954
0.3756
0.3766 0.4028
NASDAQ 1
0.2035
0.1907
0.2938
0.2141
0.6737 0.6838
3
0.2684
0.2677
0.3326
0.2879
0.3546 0.3435
7
0.3849
0.3726
0.4076
0.3762
0.4178 0.4170
Entries are Stress-period (2020) SMAPE mean (standard deviation) across 20 seeds where applicable. Deterministic GARCH benchmarks omit standard deviations.
The results in Tables.67 demonstrate that the proposed GARCH-RNN (both GARCH-GRU and GARCH-LSTM) framework maintains a clear and substantial advantage over all benchmark models during the COVID-19 stress period. Both MAE and SMAPE increase sharply across all models due to unprecedented market turbulence, yet the GARCH-RNN architectures exhibit markedly smaller error escalations than both classical GARCH specifications and non-embedded neural networks. This confirms that integrating GARCH dynamics directly into the recurrent structure provides essential stability and adaptability under extreme volatility, where conventional econometric models and purely data-driven deep learning methods struggle to accommodate abrupt regime shifts. Within the GARCH-RNN family, a distinct pattern emerges. GARCH-LSTM achieves the lowest MAE across most datasets and horizons, particularly at the one-day and three-day horizons where shocks propagate rapidly. Its more elaborate gating mechanism appears to provide additional memory depth that becomes advantageous when volatility levels remain elevated for extended periods. However, GARCH-GRU remains highly competitive, and in several settings, especially for the S&P 500, it attains the best SMAPE values, indicating stronger proportional accuracy when volatility spikes. The performance gap between the two GARCH-RNN variants is small, especially relative to the gap separating both models from the benchmarks, reinforcing the conclusion that the primary performance gains arise from the embedded GARCH–neural interaction itself rather than from the choice of recurrent cell.
Benchmark models deteriorate much more severely under stress. The two-stage hybrid (bm_GARCH_LSTM) performs considerably worse than either of the embedded GARCH-RNN models, underscoring the importance of joint end-to-end learning for capturing rapidly shifting volatility dynamics. Transformer models, despite their strong performance in benign environments, experience notable degradation, revealing limitations of purely data-driven architectures when confronted with large structural breaks. Classical GARCH and GJR-GARCH models exhibit the weakest performance overall, reflecting their inability to adapt to sudden persistence changes and nonlinear amplification common in crisis periods.
Table 8: Average Training Time per Epoch Comparison
Model Avg. Time per Epoch (s)
GARCH_GRU 665.78
GARCH_LSTM 1838.81
The reported average training times are aggregated across all forecasting horizons, datasets, and random seeds (20 runs in total).
Furthermore, to evaluate the computational efficiency, we measured the average training time per epoch for both GARCH-GRU and GARCH-LSTM models across 20 runs with different random seeds as shown in Table.8. Aggregating the results across all forecasting horizons, datasets, and seeds, the GARCH-GRU model exhibits an average epoch time of 665.78 seconds, while the GARCH-LSTM requires 1838.81 seconds. This implies that, on average, the GARCH-GRU architecture trains approximately 2.76 times faster than its LSTM counterpart. The superior efficiency of GARCH-GRU arises from its simpler gating structure, which reduces the number of recurrent parameters and matrix multiplications at each time step. Consequently, the GRU-based hybrid model delivers a substantial reduction in computational cost without sacrificing predictive accuracy, making it a more scalable and time-efficient framework for multi-horizon or high-frequency volatility forecasting applications.
Overall, these results consistently demonstrate that coupling a recurrent architecture with a GARCH component enables the model to leverage both the parametric interpretability of volatility dynamics and the nonlinear representational power of deep networks. Among all tested frameworks, the GARCH-GRU offers superior performance in stable or moderately volatile markets, while GARCH-LSTM gains a relative edge under extreme stress, yet both significantly outperform all benchmarks. These findings validate the GARCH-RNN framework as a flexible, interpretable, and empirically powerful approach to volatility forecasting across diverse market regimes.